A 100KG barrel with a raduis of 50 cm has two ropes wrapped around it , the barrel is released from rest , causing the ropes to unwind and the barrel fall spinning toward the ground . what is the speed of the barrel after it has fallen a distance of 10m ? what is the tension in each rope ? assume that the barrels mass is uniformaly distributed and that the barrel rotates as a solid cylinder >
1
Expert's answer
2020-06-22T11:12:47-0400

Apply law of conservation of energy. The moment of inertia of a solid cylinder is


"I=\frac{1}{2}mr^2."


The initial potential energy will convert into kinetic energy of translational and rotational motion:

"mgh=\frac{mv^2}{2}+\frac{I\omega^2}{2},"

where


"\omega=\frac{v}{r}."

Find the speed:


"mgh=\frac{mv^2}{2}+\frac{(mr^2/2)(v/r)^2}{2}=\frac{3}{4}mv^2,\\\space\\
v=2\sqrt{\frac{gh}{3}}=2\sqrt{\frac{9.8\cdot10}{3}}=11.43\text{ m/s}."